Jun 5, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. Eastern

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL
  4. BCAL
  5. Reconstruction and Simulation
  6. D3 update
  7. Calibration
  8. Any other business


Attendees: Manuel, Zach, Shayne, Laura, Yi, Mark, David, Elton (JLab); Zisis, Tegan (UofR); Christina, Stratos, George (Athens)

  1. Announcements
    • Users Meeting this week. Elton has also been elected as a member at large of the UGBOD.
  2. Action Items
  3. FCAL (Manuel) HDFCAL log book
    • Overnight runs completed for LED and cosmic triggers
    • Checking the stability of environmental monitors, temperature, humidity and light. See the 3 hour temperature cycle seen with BCAL
    • Investigation of individual channels that do not stabilize
    • Cosmic trigger. Overflows seen in about 1% of the triggers
    • LED and cosmic-ray data will be used to check normalization of gains. The LED data will be used to smooth out behavior between neighbors. The cosmic-ray data will be investigated for use to equalize response.
    • Zach: The FCAL has spare environmental monitoring channels that could be used to record temperature and humidity outside the darkbox. He was encouraged to work with Dave Butler to add these to the system.
    • Andrei: Asked about an apparent "reflection" in the pulse shape from a cosmic-ray. It does not seem to be a reflection, as it is not seen with LED data.
  4. BCAL HDBCAL log book
    1. Mark showed preliminary results of the cosmic-ray data. The integral and max values were plotted.
      • The inner layer shows a peak at the expected energy deposition of minimum ionizing tracks. This corresponds to about 22 MeV for the peak and 230 samples for the integral -> integral counts * 0.1 ~ MeV
    2. Elton summarized work with the LEDs
      • Data have been taken overnight with the pulser at 12, 18 and 24 degrees
      • We have learned how to operate the chiller during this time. There was a hardware setting limiting the operation of the compressor, which has been relaxed. It is now maintaining the temperature very stable, within 0.1 deg C and the 20 min period of temperature cycling is gone.
      • We ran Orlando's configuration scripts to verify operation of north and south MPPCs. All south MPPCs are operational. Likely all north MPPCs are also ok, but some software glitches need to be ironed out to confirm.
  5. Reconstruction and Simulation
    • Irina is including the actual geometry of BCAL cells into the MC
    • Andrei expects that an initial set of simulation of cosmic-ray data will be available next week. This includes energy deposition in each cell as well as attenuated energy to the end for comparison with data.
    • Tegan continues to work on the simulation program, but may need additional assistance from the offline group. Once he is up to speed, he will go through the BCAL reconstruction code with the assistance of the Reconstruction group to ensure that a) our procedures are final (e.g. time-walk is done in a single pass right now, but we may want to do it in two passes), b) locate constants that need to be placed in ccdb, c) document code as needed, b) check it into the repository or use by others (George Vassileiadis, new Regina grad student in the fall). In June he will also go over PWA and BDT reading, to be prepared for his visit to MIT in July.
  6. D3 update
    • Matt has updated sections for the FCAL and checked them in. There are still gaps that he is working on.
  7. Calibration
    • Zisis and undergraduates will work on entering calibration constants from cosmics and LED data into ccdb, with Sean's help.
  8. Any other business
    • Shift taker's documentation needs to be completed over the next week. Initial draft for BCAL is completed. Elton will meet with Manuel to go over the FCAL draft. John will be back next week and can help complete the documentation